### v2

Abstract
Due to their importance in various emerging applications, efficient algorithms for solving minimax problems have recently received increasing attention. However, many existing algorithms require prior knowledge of the problem parameters in order to achieve optimal iteration complexity. In this paper, we propose two completely parameter-free alternating gradient projection algorithms, i.e., the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm, to solve the smooth nonconvex-strongly concave and nonconvex-concave minimax problems respectively using a backtracking strategy, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concave constant $\mu$. Moreover, we show that the total number of gradient calls of the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm to obtain an $\varepsilon$-stationary point is upper bounded by $\mathcal{O}\left(L\kappa^{3}\varepsilon^{-2}\right)$ and $\mathcal{O}\left(L^{4}\varepsilon^{-4}\right)$ respectively, where $\kappa$ is the condition number. As far as we know, the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm are the first completely parameter-free algorithms for solving nonconvex-strongly concave minimax problems and nonconvex-concave minimax problems respectively. Numerical results validate the efficiency of the proposed PF-AGP algorithm.
Keywords: minimax optimization problem; backtracking strategies; parameter-free alternating gradient projection algorithm; iteration complexity; machine learning
1 Introduction
We consider the following minimax optimization problem:
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}f(x,y),$ (P)
where $\mathcal{X}\subseteq\mathbb{R}^{n}$ and $\mathcal{Y}\subseteq\mathbb{R}^{m}$ are nonempty closed and bounded convex sets, and $f:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ is a smooth function. Minimax problem is a critical class of optimization problems, known for its extensive research history and broad applications in areas such as machine learning, signal processing, and various other fields in recent years. Many practical problems can be formulated as in (P), including generative adversarial networks (GANs) (Arjovsky et al., 2017; Goodfellow et al., 2014; Sanjabi et al., 2018), reinforcement learning (Dai et al., 2018), power control and transceiver design in signal processing (Lu et al., 2020), distributed nonconvex optimization (Giannakis et al., 2017; Giordano et al., 2018; Mateos et al., 2010), robust optimization (Ben and Tal, 2009; Gao and Kleywegt, 2023), statistical learning (Giordano et al., 2018; Shafieezadeh Abadeh et al., 2015) and several other applications.
There are many research results on convex-concave minimax problems, i.e., $f(x,y)$ is convex with respect to $x$ and concave with respect to $y$. Nemirovski (Nemirovski, 2004) introduced the mirror-prox method, and Nestrov (Nesterov, 2007) proposed a dual extrapolation algorithm, both of which achieve an optimal iteration complexity of $\mathcal{O}(\varepsilon^{-1})$ for solving convex-concave minimax problems. For more details, see (Vandenberghe and Boyd, 2004; Chen et al., 2014, 2017; Dang and Lan, 2014; Monteiro and Svaiter, 2010; Nedić and Ozdaglar, 2009; Yang et al., 2020) and references therein.
Recently, various algorithms have been proposed for nonconvex-(strongly) concave minimax problems. For nonconvex-strongly concave minimax problems, several recent works (Jin et al., 2020; Lin et al., 2020a; Rafique et al., 2022; Lu et al., 2020) have proposed various algorithms that achieve a gradient complexity of $\tilde{\mathcal{O}}(\kappa_{y}^{2}\varepsilon^{-2})$ for stationary points of $\Phi(\cdot)=\max_{y\in\mathcal{Y}}f(\cdot,y)$ (when $\mathcal{X}=\mathbb{R}^{n}$), or for stationary points of $f$, where $\kappa_{y}$ is the condition number for $f(x,\cdot)$. (Lin et al., 2020b) proposed an accelerated algorithm that improves the gradient complexity bound to $\tilde{\mathcal{O}}(\sqrt{\kappa_{y}}\varepsilon^{-2})$. Furthermore, Zhang et al. (Zhang et al., 2021) proposed a generic acceleration framework which can improve the iteration complexity to $\mathcal{O}\left(\sqrt{\kappa_{y}}\varepsilon^{-2}\right)$.
For general nonconvex-concave minimax problems, there are two types of algorithms, i.e., multi-loop algorithms and single-loop algorithms. Various multi-loop algorithms have been proposed in (Kong and Monteiro, 2021; Lin et al., 2020b; Nouiehed et al., 2019; Ostrovskii et al., 2021; Rafique et al., 2022; Thekumparampil et al., 2019). The best known iteration complexity of multi-loop algorithms for solving nonconvex-concave minimax problems is $\tilde{\mathcal{O}}(\varepsilon^{-2.5})$, which was achieved by (Lin et al., 2020b). For solving nonconvex-linear minimax problems, Pan et al. (Pan et al., 2021) proposed a new alternating gradient projection algorithm with the iteration complexity of $\mathcal{O}(\varepsilon^{-3})$. Various single-loop algorithms have also been proposed, e.g., the gradient descent-ascent (GDA) algorithm (Lin et al., 2020a), the hybrid block successive approximation (HiBSA) algorithm (Lu et al., 2020), the unified single-loop alternating gradient projection (AGP) algorithm (Xu et al., 2023), and the smoothed GDA algorithm (Zhang et al., 2020). Both the AGP algorithm and the smoothed GDA algorithm achieve the best known iteration complexity, i.e., $\mathcal{O}(\varepsilon^{-4})$, among single-loop algorithms for solving nonconvex-concave minimax problems.
It should be noted that for most of the existing results mentioned above, achieving optimal complexity of the algorithm requires the assumption of knowing precise information about some parameters of the problem, such as the Lipschitz constant, or the strongly concave constant $\mu$, etc. Accurately estimating these parameters is often challenging, and conservative estimates can significantly impact algorithm performance (Lan et al., 2023). Therefore, designing parameter-free algorithms with complexity guarantees without relying on inputs of these parameters has attracted considerable attention recently.
For nonconvex-strongly concave minimax problems, there are two main parameter-free algorithms: one based on AdaGrad stepsizes and the other based on backtracking. In the framework of AdaGrad stepsizes, Yang et al. (Yang et al., 2022) proposed NeAda, which can achieve convergence rates of $\tilde{\mathcal{O}}(\varepsilon^{-2})$ and $\tilde{\mathcal{O}}(\varepsilon^{-4})$ in deterministic and stochastic settings, respectively. Li et al. (Li et al., 2022) proposed TiAda, achieving convergence rates of ${\mathcal{O}}(\varepsilon^{-2})$ and ${\mathcal{O}}(\varepsilon^{-4})$ in deterministic and stochastic settings, respectively. In the framework of backtracking, Xu et al. (Xu et al., 2024) proposed a stochastic gradient descent ascent (GDA) method with backtracking (SGDA-B), achieving an upper bound on the total number of gradient calls of $\mathcal{O}(L\kappa^{2}\log(\kappa)\varepsilon^{-2})$ with $\mathcal{O}(L\kappa^{2}\varepsilon^{-2})$ gradient calls required per backtracking step. Meanwhile, Zhang et al. (Zhang et al., 2024) proposed a proximal alternating gradient descent ascent method (AGDA+), achieving an upper bound on the total number of gradient calls of $\mathcal{O}(L\kappa^{4}\varepsilon^{-2})$ with $\mathcal{O}(1)$ gradient calls required per backtracking step. Both methods do not require knowledge of the specific value of Lipschtiz constant $L$, but still require information about the strongly concave constant $\mu$ in the function. In the framework of backtracking, Xu et al. (Xu et al., 2024) proposed a stochastic GDA method with backtracking (SGDA-B) for nonconvex-concave minimax problems, achieving an upper bound on the total number of gradient calls of $\tilde{\mathcal{O}}(L^{3}\varepsilon^{-4})$ with $\mathcal{O}(L^{3}\varepsilon^{-4})$ gradient calls required per backtracking step.
1.1 Contributions
In this paper, based on the framework of backtracking, we propose two completely parameter-free alternating gradient projection algorithms, i.e., the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm, for solving nonconvex-(strongly) concave minimax problems respectively, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concave constant $\mu$.
Moreover, we show that the total number of gradient calls of the PF-AGP-NSC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-strongly concave minimax problems is upper bounded by $\mathcal{O}\left(L\kappa^{3}\varepsilon^{-2}\right)$ where $\kappa$ is the condition number, while the total number of gradient calls of the PF-AGP-NC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-concave minimax problems is upper bounded by $\mathcal{O}\left(L^{4}\varepsilon^{-4}\right)$. As far as we know, the PF-AGP-NSC algorithm is the first completely parameter-free algorithm for solving nonconvex-strongly concave minimax problems, and the PF-AGP-NC algorithm is the first completely parameter-free algorithm which achieves the best iteration complexity in single loop algorithms for solving nonconvex-concave minimax problems.
1.2 Organization
In Sect. 2, we propose two parameter-free alternating gradient projection (PF-AGP) algorithms for nonconvex-strongly concave and nonconvex-concave minimax problems, and we then analyze their corresponding gradient complexity under two different settings. We report some numerical results in Sect. 3 and make some conclusions in the last section.
[…]
$f(y)\geq f(x)+\langle\nabla f(x),y-x\rangle+\frac{\theta}{2}\|y-x\|^{2}.$ (1)
If $-f$ satisfies (1), $f(\cdot)$ is called $\theta$-strongly concave. Denote the sign function as $\mbox{sgn}(x)$, i.e.,
$\mbox{sgn}(x)=\begin{cases}1&\text{if }x>0,\\ -1&\text{if }x\leq 0.\end{cases}$ (2)
2 Two completely parameter-free alternating gradient projection algorithms
In this section, we propose two completely parameter-free alternating gradient projection (PF-AGP) algorithms for solving the nonconvex-strongly concave and the nonconvex-concave minimax problems (P) respectively. Both algorithms do not require prior knowledge of Lipschitz constants, and the algorithm under nonconvex-strong concave settings also does not require prior knowledge of strong concave constants. The two proposed algorithms are based on the AGP algorithm (Xu et al., 2023), which uses the gradient of a regularized version of the original function, i.e.,
$f_{k}(x,y)=f(x,y)-\frac{c_{k}}{2}\|y_{k}\|^{2},$
where $c_{k}\geq 0$ is a regularization parameters. At the $k$-th iteration, AGP algorithm consists of the following two gradient projection steps for updating both $x$ and $y$:
$x_{k+1}$ $=\mathcal{P}_{\mathcal{X}}\left(x_{k}-\tfrac{1}{\beta_{k}}\nabla_{x}f(x_{k},y_{k})\right),$
$y_{k+1}$ $=\mathcal{P}_{\mathcal{Y}}\left(y_{k}+\tfrac{1}{\gamma_{k}}\nabla_{y}{f}(x_{k},y_{k})-\frac{1}{\gamma_{k}}c_{k}y_{k}\right),$
where $\mathcal{P}_{\mathcal{X}}$ and $\mathcal{P}_{\mathcal{Y}}$ is the projection operator onto $\mathcal{X}$ and $\mathcal{Y}$, respectively, and $\beta_{k}>0$, $\gamma_{k}>0$ are stepsize parameters.
Although the AGP algorithm achieves the optimal iteration complexity among single loop algorithms for nonconvex-(strongly) concave minimax problems, it requires the knowledge of the Lipschtiz constant $L$ and the strongly concave coefficient $\mu$, or at least knowledge of the Lipschtiz constant $L$ to calculate the stepsizes $\beta_{k}$ and $\gamma_{k}$ under the nonconvex-strongly concave setting and the nonconvex-concave setting, respectively, which limits its applicability. To eliminate the need for the prior knowledge of the Lipschtiz constant $L$ or the strongly concave coefficient $\mu$, we propose two parameter-free AGP algorithms by employing backtracking strategies for solving the nonconvex-strongly concave and the nonconvex-concave minimax problems, respectively. Similar backtracking strategies have been used in (Lan et al., 2023; Liu and Luo, 2022; Nesterov and Polyak, 2006; Nesterov, 2015).
2.1 Nonconvex-strongly Concave Setting
We first propose a completely parameter-free AGP algorithm for solving nonconvex-strongly concave minimax problems. More specifically, in order to estimate the Lipschitz constants for $\nabla_{x}f(\cdot,y)$, $\nabla_{y}f(\cdot,y)$, $\nabla_{y}f(x,\cdot)$ denoted as $L_{11}$, $L_{12}$ and $L_{22}$, respectively, and the strongly concave constant $\mu$, at each iteration of the proposed algorithm, we aim to find a tuple $(l_{11}^{k,i},l_{12}^{k,i},l_{22}^{k,i},\mu_{k,i})$ by backtracking such that the following conditions for $(x_{k,i},y_{k,i})$ are satisfied:
$C_{1}^{k,i}$ $=f(x_{k,i},y_{k})-f(x_{k},y_{k})-\langle\nabla_{x}f(x_{k},y_{k}),x_{k,i}-x_{k}\rangle-\frac{l^{k,i}_{11}}{2}\|x_{k,i}-x_{k}\|^{2}\leq 0,$ (C1)
$C_{2}^{k,i}$ $=\|\nabla_{y}f(x_{k,i},y_{k})-\nabla_{y}f(x_{k},y_{k})\|-l_{12}^{k,i}\|x_{k,i}-x_{k}\|\leq 0,$ (C2)
$C_{3}^{k,i}$ $=\|\nabla_{y}f(x_{k,i},y_{k,i})-\nabla_{y}f(x_{k,i},y_{k})\|-l_{22}^{k,i}\|y_{k,i}-y_{k}\|\leq 0,$ (C3)
$C_{4}^{k,i}$ $=\langle\nabla_{y}f(x_{k,i},y_{k,i})-\nabla_{y}f(x_{k,i},y_{k}),y_{k,i}-y_{k}\rangle+\mu_{k,i}\|y_{k,i}-y_{k}\|^{2}\leq 0.$ (C4)
Otherwise, we enlarge $l_{11}^{k,i},l_{12}^{k,i},l_{22}^{k,i}$ by a factor of $2$ or reduce $\mu_{k,i}$ by half according to the sign of $C_{1}^{k,i}$, $C_{2}^{k,i}$, $C_{3}^{k,i}$ and $C_{4}^{k,i}$, respectively. We then update $\beta_{k}$ and $\gamma_{k}$ by the new estimates, i.e., $l_{11}^{k,i},l_{12}^{k,i},l_{22}^{k,i},\mu_{k,i}$. The proposed algorithm for solving nonconvex-strongly concave problems, denoted as PF-AGP-NSC, is formally presented in Algorithm 1.
Algorithm 1 A parameter-free alternating gradient projection (PF-AGP-NSC) algorithm for nonconvex-strongly concave minimax problems
  Step 1: Input $x_{1},y_{1}$, $\beta_{1}$, $\gamma_{1}$, $l_{11}^{1}$, $l_{12}^{1}$, $l_{22}^{1}$, $\mu_{1}$; Set $k=1$.
  Step 2: Updata $x_{k}$ and $y_{k}$:
   (a): Set $i=1$, $l_{11}^{k,i}=l_{11}^{k}$, $l_{12}^{k,i}=l_{12}^{k}$, $l_{22}^{k,i}=l_{22}^{k}$, $\mu_{k,i}=\mu_{k}$, $\beta_{k,i}=\beta_{k}$, $\gamma_{k,i}=\gamma_{k}$.
   (b): Update $x_{{k,i}}$ and $y_{{k,i}}$:
$x_{{k,i}}$ $=\mathcal{P}_{\mathcal{X}}\left(x_{k}-\frac{1}{\beta_{k,i}}\nabla_{x}f(x_{k},y_{k})\right),$ (3)
$y_{{k,i}}$ $=\mathcal{P}_{\mathcal{Y}}\left(y_{k}+\frac{1}{\gamma_{k,i}}\nabla_{y}f(x_{{k,i}},y_{k})\right).$ (4)
   (c): Compute $C_{1}^{k,i}$, $C_{2}^{k,i}$, $C_{3}^{k,i}$, $C_{4}^{k,i}$ as in (C1), (C2), (C3), (C4), respectively;
   (d): Update $l^{k,i}_{11}$, $l^{k,i}_{12}$, $l^{k,i}_{22}$, $\mu_{k,i}$:
$l^{k,i+1}_{11}$ $=\frac{\mbox{sgn}(C_{1}^{k,i})+3}{2}l^{k,i}_{11},\quad l^{k,i+1}_{12}=\frac{\mbox{sgn}(C_{2}^{k,i})+3}{2}l^{k,i}_{12},$ (5)
$l^{k,i+1}_{22}$ $=\frac{\mbox{sgn}(C_{3}^{k,i})+3}{2}l^{k,i}_{22},\quad\mu_{k,i+1}=\frac{2}{\mbox{sgn}(C_{4}^{k,i})+3}\mu_{k,i}.$ (6)
   (e): If $C_{1}^{k,i}\leq 0$, $C_{2}^{k,i}\leq 0$, $C_{3}^{k,i}\leq 0$ and $C_{4}^{k,i}\leq 0$, then
     $x_{k+1}=x_{k,i}$, $y_{k+1}=y_{k,i}$, $l_{11}^{k}=l_{11}^{k,i+1}$, $l_{12}^{k}=l_{12}^{k,i+1}$, $l_{22}^{k}=l_{22}^{k,i+1}$, $\mu_{k}=\mu_{k,i+1}$,
     $\beta_{k}=\beta_{k,i}$, $\gamma_{k}=\gamma_{k,i}$, go to Step 3;
     Otherwise, $i=i+1$,
$\beta_{k,i}={l_{11}^{k,i}}+\frac{l_{12}^{k,i}\mu_{k-1,i}}{8(l_{22}^{k-1,i})^{2}}+\frac{32(l_{12}^{k,i})^{2}(l_{22}^{k-1,i})^{2}}{\mu_{k,i}\mu_{k-1,i}^{2}},\quad\gamma_{k,i}=\frac{8(l_{22}^{k,i})^{2}}{\mu_{k,i}},$
     go to Step 2(b).
  Step 3: If some stationary condition is satisfied, stop; Otherwise, set $k=k+1$, go to Step 2.
It is worth noting that the parameter-free NeAda (Yang et al., 2022) and TiAda (Li et al., 2022) algorithms with adaptive step size are only used to solve the minimax problems when $x$ is unconstrained, while the proposed PF-AGP-NSC algorithm can solve the constrained case. Compared with the SGDA-B (Xu et al., 2023) algorithm and AGDA+ (Zhang et al., 2024) algorithm using the backtracking framework under the nonconvex-strongly concave setting, AGDA+ and SGDA-B still require knowledge of the strongly concave constant $\mu$, while the proposed PF-AGP-NSC algorithm is a completely parameter-free algorithm that does not require any prior knowledge of the parameters.
2.1.1 Complexity analysis
In this subsection, we analyze the iteration complexity of Algorithm 1 for solving nonconvex-strongly concave minimax optimization problems (P), i.e., $f(x,y)$ is nonconvex with respect to $x$ for any fixed $y\in\mathcal{Y}$, and $\mu$-strongly concave with respect to $y$ for any given $x\in\mathcal{X}$. We first need to make the following assumption about the smoothness of $f(x,y)$.
Assumption 1
$f(x,y)$ has Lipschitz continuous gradients, i.e., there exist positive scalars $L_{11},L_{12},L_{22}$ such that for any $x,\tilde{x}\in\mathcal{X},y,\tilde{y}\in\mathcal{Y}$,
$\|\nabla_{x}f(x,y)-\nabla_{x}f(\tilde{x},y)\|$ $\leq L_{11}\|x-\tilde{x}\|,$
$\|\nabla_{y}f(x,y)-\nabla_{y}f(\tilde{x},y)\|$ $\leq L_{12}\|x-\tilde{x}\|,$
$\|\nabla_{y}f(x,\tilde{y})-\nabla_{y}f(x,y)\|$ $\leq L_{22}\|\tilde{y}-y\|.$
We denote ${L}=\max\{L_{11},L_{12},L_{22}\}$. To analyze the convergence of Algorithm 1, we define the stationarity gap as the termination criterion as follows.
Definition 1
At each iteration of Algorithm 1, the stationarity gap for problem (P) with respect to $f(x,y)$ is defined as:
$\nabla G(x_{k},y_{k})=\begin{bmatrix}\beta_{k}\left(x_{k}-\mathcal{P}_{\mathcal{X}}\left(x_{k}-\frac{1}{\beta_{k}}\nabla_{x}f(x_{k},y_{k})\right)\right)\\ \gamma_{k}\left(y_{k}-\mathcal{P}_{\mathcal{Y}}\left(y_{k}+\frac{1}{\gamma_{k}}\nabla_{y}f(x_{k},y_{k})\right)\right)\end{bmatrix}.$ (7)
We denote $\nabla{G}_{k}=\nabla{G}(x_{k},y_{k})$, $(\nabla G_{k})_{x}=\beta_{k}(x_{k}-\operatorname{\mathcal{P}_{\mathcal{X}}}(x_{k}-\frac{1}{\beta_{k}}\nabla_{x}f(x_{k},y_{k})))$, and $(\nabla G_{k})_{y}=\gamma_{k}(y_{k}-\operatorname{\mathcal{P}_{\mathcal{Y}}}(y_{k}+\frac{1}{\gamma_{k}}\nabla_{y}f(x_{k},y_{k})))$.
Definition 2
At each iteration of Algorithm 1, the stationarity gap for problem (P) with respect to $f_{k}(x,y)$ is defined as:
$\nabla\tilde{G}(x_{k},y_{k})=\begin{bmatrix}\beta_{k}\left(x_{k}-P_{x}\left(x_{k}-\frac{1}{\beta_{k}}\nabla_{x}f_{k}(x_{k},y_{k})\right)\right)\\ \gamma_{k}\left(y_{k}-P_{y}\left(y_{k}+\frac{1}{\gamma_{k}}\nabla_{y}f_{k}(x_{k},y_{k})\right)\right)\end{bmatrix}.$
Under Assumption 1, we first establish the following lemma, which provides bounds on the changes in the function value when $x_{k}$ is updated at each iteration of Algorithm 1. This proof is similar to that of Lemma 2.1 in (Xu et al., 2023), for the sake of completeness, we give its proof.
Lemma 3
Suppose that Assumption 1 holds. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, then we have
$f(x_{k+1},y_{k})-f(x_{k},y_{k})\leq-\left(\beta_{k}-\frac{l^{k}_{11}}{2}\right)\|x_{k+1}-x_{k}\|^{2}.$ (8)
Proof  By the optimality condition for (3), we have
$\langle\nabla_{x}f\left(x_{k},y_{k}\right)+\beta_{k}\left(x_{k+1}-x_{k}\right),x_{k}-x_{k+1}\rangle\geq 0.$ (9)
The backtracking strategy (C1) implies that
$f\left(x_{k+1},y_{k}\right)-f\left(x_{k},y_{k}\right)\leq\langle\nabla_{x}f\left(x_{k},y_{k}\right),x_{k+1}-x_{k}\rangle+\tfrac{l^{k}_{11}}{2}\|x_{k+1}-x_{k}\|^{2}.$ (10)
By adding (9) and (10), we obtain
$f\left(x_{k+1},y_{k}\right)-f\left(x_{k},y_{k}\right)$ $\leq-\left(\beta_{k}-\tfrac{l^{k}_{11}}{2}\right)\|x_{k+1}-x_{k}\|^{2},$
which completes the proof.
Next, we provide an estimate of upper bound of the difference between $f(x_{k+1},y_{k+1})$ and $f(x_{k},y_{k})$. First, we need to make the following assumption on the parameter $\gamma_{k}$.
Assumption 2
$\{\gamma_{k}\}$ is a nonnegative monotonically increasing sequence.
Lemma 4
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, then we have
$f(x_{k+1},y_{k+1})-f(x_{k},y_{k})$
$\quad\leq-\left(\beta_{k}-\frac{l_{11}^{k}}{2}-\frac{l_{12}^{k}}{2\gamma_{k-1}}\right)\|x_{k+1}-x_{k}\|^{2}+{\gamma_{k-1}}\|y_{k+1}-y_{k}\|^{2}$
$\quad\quad-\left(\mu_{k-1}-\frac{\gamma_{k-1}}{2}-\frac{(l_{22}^{k-1})^{2}}{2\gamma_{k-1}}\right)\|y_{k}-y_{k-1}\|^{2}.$ (11)
Proof  The optimality condition for $y_{k}$ in (4) implies that
$\langle\nabla_{y}f(x_{k},y_{k-1})-\gamma_{k-1}(y_{k}-y_{k-1}),y_{k+1}-y_{k}\rangle\leq 0.$ (12)
By the concavity of $f(x,y)$ with respect to $y$, and combining (12), we have
$f(x_{k+1},y_{k+1})-f(x_{k+1},y_{k})$
$\leq$ $\langle\nabla_{y}f(x_{k+1},y_{k}),y_{k+1}-y_{k}\rangle$
$\leq$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle+\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle.$ (13)
Denoting $v_{k+1}=(y_{k+1}-y_{k})-(y_{k}-y_{k-1})$, we can write the first inner product term in the r.h.s. of (13) as
$\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle$
$=$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle$
$+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle.$ (14)
Next, we estimate the three terms on the right-hand side of (14) respectively. By the backtracking strategy (C2), (C3) and the Cauchy-Schwarz inequality, we have
$\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle\leq\frac{l_{12}^{k}}{2\gamma_{k}}\|x_{k+1}-x_{k}\|^{2}+\frac{\gamma_{k}}{2}\|y_{k+1}-y_{k}\|^{2}$ (15)
and
$\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle\leq\frac{(l_{22}^{k-1})^{2}}{2\gamma_{k-1}}\|y_{k}-y_{k-1}\|^{2}+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}.$ (16)
By the backtracking strategy (C4) of $\mu_{k-1}$, we obtain
$\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle\leq-\mu_{k-1}\|y_{k}-y_{k-1}\|^{2}.$ (17)
Moreover, it can be easily checked that
$\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle=\frac{1}{2}\|y_{k}-y_{k-1}\|^{2}+\frac{1}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{1}{2}\|v_{k+1}\|^{2}.$ (18)
Plugging (14)-(18) into (13) and rearranging the terms, we conclude that
$f(x_{k+1},y_{k+1})-f(x_{k+1},y_{k})$ $\leq\frac{l_{12}^{k}}{2\gamma_{k}}\|x_{k+1}-x_{k}\|^{2}-\left(\mu_{k-1}-\frac{\gamma_{k-1}}{2}-\frac{(l_{22}^{k-1})^{2}}{2\gamma_{k-1}}\right)\|y_{k}-y_{k-1}\|^{2}$
$\quad+\left(\frac{\gamma_{k}}{2}+\frac{\gamma_{k-1}}{2}\right)\|y_{k+1}-y_{k}\|^{2}.$ (19)
By Lemma 8, we have
$f(x_{k+1},y_{k})-f(x_{k},y_{k})\leq-(\beta_{k}-\frac{l_{11}^{k}}{2})\|x_{k+1}-x_{k}\|^{2}.$ (20)
The proof is then completed by combining (19) with (20).
We now establish an important recursion for Algorithm 1.
Lemma 5
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, and denoting
$f_{k+1}$ $=f(x_{k+1},y_{k+1}),\quad S_{k+1}=\frac{2\gamma^{2}_{k}}{\mu_{k}}\|y_{k+1}-y_{k}\|^{2},$
$F_{k+1}$ $=f_{k+1}+S_{k+1}-\left(\mu_{k}+\frac{7\gamma_{k}}{2}-\frac{(l^{k}_{22})^{2}}{2\gamma_{k}}-\frac{2(l^{k}_{22})^{2}}{\mu_{k}}\right)\|y_{k+1}-y_{k}\|^{2}-\frac{2\gamma_{k}^{2}}{\mu_{k}}\mathcal{D}^{2}_{y},$ (21)
then we have
$F_{k+1}-F_{k}\leq$ $-\left(\frac{2\mu_{k}\gamma_{k}-(l_{22}^{k})^{2}}{2\gamma_{k}}+\frac{\mu_{k}\gamma_{k}-4(l_{22}^{k})^{2}}{2\mu_{k}}\right)\|y_{k+1}-y_{k}\|^{2}$
$\quad-\left(\beta_{k}-\frac{l_{11}^{k}}{2}-\frac{l_{12}^{k}}{2\gamma_{k-1}}-\frac{2(l_{12}^{k})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}\right)\|x_{k+1}-x_{k}\|^{2},$ (22)
where $\mathcal{D}_{y}=\max\{\|y_{1}-y_{2}\|\mid y_{1},y_{2}\in\mathcal{Y}\}$.
Proof  By the optimality condition for $y_{k}$ in (4) we obtain
$\langle\nabla_{y}f(x_{k+1},y_{k})-\gamma_{k}(y_{k+1}-y_{k}),y_{k}-y_{k+1}\rangle\leq 0.$ (23)
Combining (23) and (12), we have
$\langle\nabla_{y}f(x_{k},y_{k-1})-\nabla_{y}f(x_{k+1},y_{k})+\gamma_{k}(y_{k+1}-y_{k})-\gamma_{k-1}(y_{k}-y_{k-1}),y_{k+1}-y_{k}\rangle\leq 0,$ (24)
which implies that
$\gamma_{k}\|y_{k+1}-y_{k}\|^{2}-\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle\leq\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle.$ (25)
Denoting $v_{k+1}=(y_{k+1}-y_{k})-(y_{k}-y_{k-1})$, and using the identity $\frac{1}{2}\langle v_{k+1},y_{k+1}-y_{k}\rangle=\frac{1}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{1}{2}\|v_{k+1}\|^{2}-\frac{1}{2}\|y_{k}-y_{k-1}\|^{2}$, we conclude that
$\gamma_{k}\|y_{k+1}-y_{k}\|^{2}-\gamma_{k-1}\left(\frac{1}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{1}{2}\|y_{k}-y_{k-1}\|^{2}-\frac{1}{2}\|v_{k+1}\|^{2}\right)$
$\quad\leq\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle.$ (26)
By rearranging the terms of the above inequality, we then have
$\left(\gamma_{k}-\frac{\gamma_{k-1}}{2}\right)\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}-\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$
$\leq$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle$
$=$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),v_{k+1}\rangle$
$+\langle\nabla_{y}f(x_{k},y_{k})-\nabla_{y}f(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle.$ (27)
By backtracking strategy (C2) and the Cauchy-Schwarz inequality, we can easily get that
$\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle\leq\frac{(l_{12}^{k})^{2}}{2\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+\frac{\mu_{k}}{2}\|y_{k+1}-y_{k}\|^{2}.$ (28)
Combining (15), (16), (27) and (28), we have
$\frac{\gamma_{k}}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$ $\leq\frac{(l_{12}^{k})^{2}}{2\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+\left(\frac{\mu_{k}}{2}+\frac{\gamma_{k-1}}{2}-\frac{\gamma_{k}}{2}\right)\|y_{k+1}-y_{k}\|^{2}$
$\quad-\left(\mu_{k-1}-\frac{(l_{22}^{k-1})^{2}}{2\gamma_{k-1}}\right)\|y_{k}-y_{k-1}\|^{2}.$ (29)
By multiplying $\frac{4\gamma_{k-1}}{\mu_{k-1}}$ on both sides of (29), using the definition of $S_{k+1}$ and Assumption 2, we obtain
$\begin{split}S_{k+1}-S_{k}&\leq\frac{2(l_{12}^{k})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}\|x_{k+1}-x_{k}\|^{2}+2\gamma_{k-1}\|y_{k+1}-y_{k}\|^{2}\\ &\quad-\left({4\gamma_{k-1}}-\frac{2(l_{22}^{k-1})^{2}}{\mu_{k-1}}\right)\|y_{k}-y_{k-1}\|^{2}+\left(\frac{2\gamma_{k}^{2}}{\mu_{k}}-\frac{2\gamma^{2}_{k-1}}{\mu_{k-1}}\right)\mathcal{D}_{y}^{2}.\end{split}$
The proof is then completed by Lemma 4 and the definition.
We are now ready to establish the iteration complexity for the PF-AGP-NSC algorithm. In particular, letting $\nabla G(x_{k},y_{k})$ be defined as in Definition 1 and $\varepsilon>0$ be a given target accuracy, we provide a bound on $T(\varepsilon)$, the first iteration index to achieve an $\varepsilon$-stationary point, i.e., $\|\nabla G(x_{k},y_{k})\|\leq\varepsilon$, which is equivalent to
$T(\varepsilon)=\min\{k\mid\|\nabla G(x_{k},y_{k})\|\leq\varepsilon\}.$ (30)
Theorem 6
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 1, $\beta_{k}={l_{11}^{k}}+\frac{l_{12}^{k}\mu_{k-1}}{8(l_{22}^{k-1})^{2}}+\frac{32(l_{12}^{k})^{2}(l_{22}^{k-1})^{2}}{\mu_{k}\mu_{k-1}^{2}}$, $\gamma_{k}=\frac{8(l_{22}^{k})^{2}}{\mu_{k}}$, then it holds that
${T}(\varepsilon)\leq\frac{F_{2}-\underline{F}}{d_{1}\varepsilon^{2}}+1,$ (31)
where $F_{2}$ is defined in (21), $\underline{F}=\underline{f}-\left(\frac{224L_{22}^{2}}{\mu}+\mu_{1}-\frac{2(l_{22}^{1})^{2}}{\mu_{1}}-\frac{\left(l_{22}^{1}\mu\right)^{2}}{128L_{22}^{2}}+\frac{2^{13}L_{22}^{4}}{\mu^{3}}\right)\mathcal{D}^{2}_{y}$, with $\underline{f}=\min_{(x,y)\in X\times Y}f(x,y)$, $\mathcal{D}_{y}=\max\{\|y_{1}-y_{2}\|\mid y_{1},y_{2}\in\mathcal{Y}\}$, $d_{1}=\min\{\frac{\beta_{\text{min}}}{2\beta_{\text{max}}^{2}+4L^{2}_{12}},\frac{15\mu^{3}+2^{7}l_{22}^{1}\mu}{2^{15}L_{22}^{4}}\}$, $\beta_{\text{min}}={l^{1}_{11}}+\frac{l^{1}_{12}\mu}{64(L_{22})^{2}}+\frac{32(l_{12}^{1})^{2}(l^{1}_{22})^{2}}{\mu_{1}^{3}}$ and $\beta_{\text{max}}=2L_{11}+\frac{\mu_{1}L_{12}}{4(l^{1}_{22})^{2}}+\frac{4096(L_{12})^{2}(L_{22})^{2}}{\mu^{3}}$.
Proof  It follows immediately from Definition 1 and (3) that
$\|(\nabla G_{k})_{x}\|=\beta_{k}\|x_{k+1}-x_{k}\|.$ (32)
On the other hand, by the triangle inequality, we obtain that
$\|(\nabla G_{k})_{y}\|$ $\leq\gamma_{k}\left\|P_{y}\left(y_{k}+\frac{1}{\gamma_{k}}\nabla_{y}f(x_{k+1},y_{k})\right)-P_{y}\left(y_{k}+\frac{1}{\gamma_{k}}\nabla_{y}f(x_{k},y_{k})\right)\right\|$
$\quad+\gamma_{k}\|y_{k+1}-y_{k}\|$
$\leq L_{12}\|x_{k+1}-x_{k}\|+\gamma_{k}\|y_{k+1}-y_{k}\|.$ (33)
By combining (32) and (33), and using the Cauchy-Schwarz inequality, we obtain
$\|\nabla G_{k}\|^{2}$ $\leq(\beta_{k}^{2}+2L_{12}^{2})\|x_{k+1}-x_{k}\|^{2}+2\gamma_{k}^{2}\|y_{k+1}-y_{k}\|^{2}.$ (34)
Let $d^{(k)}_{1}={\min\left\{\frac{\frac{2\mu_{k}\gamma_{k}-\left(l^{k}_{22}\right)^{2}}{2\gamma_{k}}+\frac{\mu_{k}\gamma_{k}-4\left(l_{22}^{k}\right)^{2}}{2\mu_{k}}}{2\gamma_{k}^{2}},\frac{\beta_{k}-\frac{l^{k}_{11}}{2}-\frac{l^{k}_{12}}{2\gamma_{k-1}}-\frac{2(l^{k}_{12})^{2}\gamma_{k-1}}{\mu_{k-1}\mu_{k}}}{\beta_{k}^{2}+2L_{12}^{2}}\right\}}>0$. Multiplying both sides of (34) by $d^{(k)}_{1}$, and using (22) in Lemma 5, we have
$d^{(k)}_{1}\|\nabla G_{k}\|^{2}$ $\leq F_{k}-F_{k+1}.$ (35)
By using the definitions of $d_{1}$, $\beta_{k}$ and $\gamma_{k}$, we have $d_{1}<d^{(k)}_{1}$, then we get
$d_{1}\|\nabla G_{k}\|^{2}\leq F_{k}-F_{k+1}.$ (36)
Summing up the above inequalities from $k=2$ to $k=T(\varepsilon)$, we obtain
$\sum_{k=2}^{T(\varepsilon)}d_{1}\|\nabla G_{k}\|^{2}$ $\leq F_{2}-F_{T(\varepsilon)+1}.$ (37)
Note that by the definition of $F_{k+1}$ in Lemma 5, we have
$F_{T(\varepsilon)+1}$ $=f_{T(\varepsilon)+1}+S_{T(\varepsilon)+1}-\frac{2\gamma_{T(\varepsilon)}^{2}}{\mu_{T(\varepsilon)}}\mathcal{D}^{2}_{y}$
$\quad-\left(\frac{7}{2}\gamma_{T(\varepsilon)+1}+\mu_{T(\varepsilon)+1}-\frac{2(l_{22}^{T(\varepsilon)+1})^{2}}{\mu_{T(\varepsilon)+1}}-\frac{(l_{22}^{T(\varepsilon)+1})^{2}}{2\gamma_{T(\varepsilon)+1}}\right)||y_{T(\varepsilon)+1}-y_{T(\varepsilon)}||^{2}$
$\geq\underline{f}-\left(\frac{224L_{22}^{2}}{\mu}+\mu_{1}-\frac{2(l_{22}^{1})^{2}}{\mu_{1}}-\frac{\left(l_{22}^{1}\mu\right)^{2}}{128L_{22}^{2}}+\frac{2^{13}L_{22}^{4}}{\mu^{3}}\right)\mathcal{D}^{2}_{y}$
$=\underline{F},$ (38)
where the inequality follows from the definitions of $\underline{f}$ and $\mathcal{D}_{y}$, and the fact that $S_{k}\geq 0$ ($\forall k\geq 1$). We then conclude from (37) that
$\sum_{k=2}^{T(\varepsilon)}d_{1}\|\nabla G_{k}\|^{2}$ $\leq{F_{2}}-F_{T(\varepsilon)+1}\leq{F}_{2}-\underline{F},$ (39)
which, in view of the definition of $T(\varepsilon)$, implies that
$d_{1}(T(\varepsilon)-1)\varepsilon^{2}\leq F_{2}-\underline{F},$ (40)
or equivalently,
$T(\varepsilon)\leq\frac{F_{2}-\underline{F}}{d_{1}\varepsilon^{2}}+1.$ (41)
We complete the proof.
Remark 7
At the $k$-th iteration of Algorithm 1, the number of backtracking step is upper bounded by $\log_{2}\left(\frac{2L_{11}}{l_{11}^{1}}\right)+\log_{2}\left(\frac{2L_{12}}{l_{12}^{1}}\right)+\log_{2}\left(\frac{2L_{22}}{l_{22}^{1}}\right)+\log_{2}\left(\frac{2\mu_{1}}{\mu}\right)=\mathcal{O}(1)$. By Theorem 6, the total number of gradient calls of Algorithm 1 to obtain an $\varepsilon$-stationary point that satisfies $\|\nabla G(x_{k},y_{k})\|\leq\varepsilon$ is upper bounded by $\mathcal{O}(L\kappa^{3}\varepsilon^{-2})$ under the nonconvex-strongly concave setting, where $\kappa$ is the condition number.
2.2 Nonconvex-concave Setting
In this subsection, we propose a completely parameter-free AGP algorithm for solving nonconvex-concave minimax problems. Similar to the idea of the PF-AGP-NSC Algorithm, we estimate the Lipschitz constants $L_{11}$, $L_{12}$ and $L_{22}$ iteratively, however we do not need to estimate the strongly concave constant $\mu$. At each iteration of the proposed algorithm, we aim to find a tuple $(l^{k,i}_{11},l^{k,i}_{12},l^{k,i}_{22})$ by backtracking such that conditions (C1), (C2), (C3) and the following condition (C5) are satisfied at $(x_{k,i},y_{k,i})$:
$C_{5}^{k,i}$ $=\langle\nabla_{y}f_{k,i}(x_{k,i},y_{k,i})-\nabla_{y}f_{k,i}(x_{k,i},y_{k}),y_{k,i}-y_{k}\rangle+\frac{c^{2}_{k,i}+c_{k,i}{l}^{k,i}_{22}}{{l}^{k,i}_{22}+2c_{k,i}}\|y_{k,i}-y_{k}\|^{2}$
$\qquad+\frac{1}{{l}^{k,i}_{22}+2c_{k,i}}\|\nabla_{y}f_{k,i}(x_{k,i},y_{k,i})-\nabla_{y}f_{k,i}(x_{k,i},y_{k})\|^{2}\leq 0.$ (C5)
The proposed algorithm, denoted as PF-AGP-NC, is formally presented in Algorithm 2.
Please note that the existing SGDA-B algorithm (Xu et al., 2023) for solving nonconvex-concave minimax problems requires the realization of the diameter of the feasible set $\mathcal{Y}$ in order to construct the key auxiliary problem, while the PF-AGP-NC algorithm we propose is a completely parameter-free algorithm that does not require any prior parameters.
Algorithm 2 A parameter-free alternating gradient projection (PF-AGP-NC) algorithm for nonconvex-concave minimax problems
  Step 1: Input $x_{1},y_{1}$, $\beta_{1}$, $\gamma_{1}$, $c_{1}$, $l_{11}^{1}$, $l_{12}^{1}$, $l_{22}^{1}$; Set $k=1$.
  Step 2: Updata $x_{k}$ and $y_{k}$:
   (a): Set $i=1$, $l_{11}^{k,i}=l_{11}^{k}$, $l_{12}^{k,i}=l_{12}^{k}$, $l_{22}^{k,i}=l_{22}^{k}$, $\beta_{k,i}=\beta_{k}$, $\gamma_{k,i}=\gamma_{k}$, $c_{k,i}=c_{k}$.
   (b): Update $x_{{k,i}}$ and $y_{{k,i}}$:
$x_{{k,i}}$ $=\mathcal{P}_{\mathcal{X}}\left(x_{k}-\frac{1}{\beta_{k,i}}\nabla_{x}f(x_{k},y_{k})\right),$ (42)
$y_{{k,i}}$ $=\mathcal{P}_{\mathcal{Y}}\left(y_{k}+\frac{1}{\gamma_{k,i}}\nabla_{y}f(x_{{k,i}},y_{k})-\frac{1}{\gamma_{k,i}}{c_{k,i}}y_{k}\right).$ (43)
   (c): Compute $C_{1}^{k,i}$, $C_{2}^{k,i}$, $C_{3}^{k,i}$, $C_{5}^{k,i}$ as in (C1), (C2), (C3), (C5);
   (d): Update $l^{k,i}_{11}$, $l^{k,i}_{12}$, $l^{k,i}_{22}$:
$l^{k,i+1}_{11}$ $=\frac{\mbox{sgn}(C_{1}^{k,i})+3}{2}l^{k,i}_{11},\quad l^{k,i+1}_{12}=\frac{\mbox{sgn}(C_{2}^{k,i})+3}{2}l^{k,i}_{12},$ (44)
$l^{k,i+1}_{22}$ $=\frac{\max\{\mbox{sgn}(C_{3}^{k,i}),\mbox{sgn}(C_{5}^{k,i})\}+3}{2}l^{k,i}_{22}.$ (45)
   (e): If $C_{1}^{k,i}\leq 0$, $C_{2}^{k,i}\leq 0$, $C_{3}^{k,i}\leq 0$ and $C_{5}^{k,i}\leq 0$, then
     $x_{k+1}=x_{k,i}$, $y_{k+1}=y_{k,i}$, $l_{11}^{k}=l_{11}^{k,i+1}$, $l_{12}^{k}=l_{12}^{k,i+1}$, $l_{22}^{k}=l_{22}^{k,i+1}$, $\beta_{k}=\beta_{k,i}$, $\gamma_{k}=\gamma_{k,i}$,
     $c_{k}=c_{k,i}$, go to Step 3;
     Otherwise, $i=i+1$,
$\beta_{k,i}=\frac{l_{12}^{k,i}}{20l^{k-1,i}_{22}}+\frac{1280(l_{12}^{k,i})^{2}\sqrt{k}}{19^{2}l_{12}^{k-1,i}},\quad\gamma_{k,i}=20l^{k,i}_{22},\quad c_{k,i}=\frac{19l^{k,i}_{22}}{k^{1/4}},$ (46)
     go to Step 2(b).
  Step 3: If some stationary condition is satisfied, stop; Otherwise, set $k=k+1$, go to Step 2.
2.2.1 Complexity Analysis
In this subsection, we analyze the iteration complexity of Algorithm 2 for solving (P) under the general nonconvex-concave setting. The backtracking strategy (C5) is a key inequality that we will use to establish some important recursions for the PF-AGP-NC algorithm under the nonconvex-concave setting in the following two lemmas. This is also one of the main differences between proofs in the nonconvex- strongly concave setting and proofs in the nonconvex-concave setting.
Lemma 8
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 2. If $\forall k$, $\beta_{k}>l^{k}_{11}$ and $\gamma_{k}\geq\frac{{l}_{22}^{k}+2c_{k}}{2}$, then
$f({x_{k+1}},y_{k+1})-f({x_{k}},y_{k})$
$\leq$ $-\left(\frac{{\beta}_{k}}{2}-\frac{\left({l_{12}^{k}}\right)^{2}}{2\gamma_{k-1}}\right)\|x_{k+1}-x_{k}\|^{2}+{\gamma_{k-1}}\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$
$\quad+\frac{c_{k-1}}{2}\left(\|y_{k+1}\|^{2}-\|y_{k}\|^{2}\right)+\frac{c_{k}}{2}\|y_{k+1}-y_{k}\|^{2}.$ (47)
Proof  The optimality condition for $y_{k}$ in (43) implies that
$\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\gamma_{k-1}(y_{k}-y_{k-1}),y_{k}-y_{k+1}\rangle\leq 0.$ (48)
The concavity of $f_{k}(x_{k+1},y)$ with respect to $y$ together with (48) then imply that
$f_{k}(x_{k+1},y_{k+1})-f_{k}(x_{k+1},y_{k})$ $\leq\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
$\quad+\langle\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle$
$=\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
$\quad+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),v_{k+1}\rangle$
$\quad+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle$
$\quad+\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle,$ (49)
where $v_{k+1}=y_{k+1}-y_{k}-(y_{k}-y_{k-1})$. We now provide upper bounds for the inner product terms on the right hand side of (49). Firstly, by the definition of $f_{k}(x_{k+1},y_{k})$ and $f_{k-1}(x_{k},y_{k})$, the backtracking strategy (C2) and the Cauchy-Schwarz inequality, we have
$\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
$=\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle-(c_{k}-c_{k-1})\langle y_{k},y_{k+1}-y_{k}\rangle$
$\leq\frac{(l_{12}^{k})^{2}}{2\gamma_{k-1}}\|x_{k+1}-x_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{c_{k}-c_{k-1}}{2}(\|y_{k+1}\|^{2}-\|y_{k}\|^{2})$
$\quad+\frac{c_{k}-c_{k-1}}{2}\|y_{k+1}-y_{k}\|^{2}$
$\leq\frac{(l_{12}^{k})^{2}}{2\gamma_{k-1}}\|x_{k+1}-x_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{c_{k}-c_{k-1}}{2}(\|y_{k+1}\|^{2}-\|y_{k}\|^{2})$
$\quad+\frac{c_{k}}{2}\|y_{k+1}-y_{k}\|^{2}.$ (50)
Secondly, by the Cauchy-Schwarz inequality, we get
$\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),v_{k+1}\rangle$
$\leq$ $\frac{1}{2\gamma_{k-1}}\|\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1})\|^{2}+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}.$ (51)
Thirdly, by (C4) we have that
$\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle$
$\leq$ $-\frac{1}{{l}_{22}^{k-1}+2c_{k-1}}\|\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1})\|^{2}$
$\quad-\frac{c_{k-1}{l}_{22}^{k-1}+c^{2}_{k-1}}{{l}_{22}^{k-1}+2c_{k-1}}\|y_{k}-y_{k-1}\|^{2}$
$\leq$ $-\frac{1}{{l}_{22}^{k-1}+2c_{k-1}}\|\nabla_{y}f_{k}(x_{k},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1})\|^{2}.$ (52)
We also have the following equation that
$\begin{split}\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle=\frac{\gamma_{k-1}}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}-\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}.\end{split}$ (53)
By $\gamma_{k}\geq\frac{{l}_{22}^{k}+2c_{k}}{2}$, plugging (50)-(53) into (49), and using the definition of $f_{k}(x_{k+1},y_{k+1})$ and $f_{k}(x_{k+1},y_{k})$, we obtain
$f(x_{k+1},y_{k+1})-f(x_{k+1},y_{k})$
$\leq$ $\frac{(l_{12}^{k})^{2}}{2\gamma_{k-1}}\|x_{k+1}-x_{k}\|^{2}+\gamma_{k-1}\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$
$\quad+\frac{c_{k-1}}{2}\left(\|y_{k+1}\|^{2}-\|y_{k}\|^{2}\right)+\frac{c_{k}}{2}\|y_{k+1}-y_{k}\|^{2}.$ (54)
By $\beta_{k}>l^{k}_{11}$, we have
$f(x_{k+1},y_{k})-f(x_{k},y_{k})\leq-\frac{\beta_{k}}{2}\|x_{k+1}-x_{k}\|^{2}.$ (55)
By adding (54) and (55), we complete the proof.
Lemma 9
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 2. Denote
$S_{k+1}$ $=\frac{8\gamma_{k}^{2}}{c_{k}}\|y_{k+1}-y_{k}\|^{2}+{8}{\gamma_{k}}\left(\frac{c_{k+1}}{c_{k}}-1\right)\|y_{k+1}\|^{2},$ (56)
$F_{k+1}$ $={f(x_{k+1},y_{k+1})}+S_{k+1}-\frac{15\gamma_{k}}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{c_{k}}{2}\|y_{k+1}\|^{2}.$ (57)
If
$\beta_{k}>l^{k}_{11},\quad\frac{\gamma_{k+1}}{c_{k+1}}-\frac{\gamma_{k}}{c_{k}}\leq\frac{1}{5},\quad\gamma_{k}\geq\frac{{l}^{k}_{22}+2c_{k}}{2},$ (58)
then $\forall k\geq 2$,
$F_{k+1}-F_{k}=$ $-\left(\frac{9\gamma_{k}}{10}-\frac{c_{k}}{2}\right)\|y_{k+1}-y_{k}\|^{2}$
$-\left(+\frac{\beta_{k}}{2}-\frac{(l^{k}_{12})^{2}}{2\gamma_{k-1}}-\frac{16(l^{k}_{12})^{2}\gamma_{k-1}}{c_{k-1}^{2}}\right)\|x_{k+1}-x_{k}\|^{2}$
$+8\left({\gamma_{k}}\frac{c_{k+1}}{c_{k}}-{\gamma_{k-1}}\frac{c_{k}}{c_{k-1}}\right)\|y_{k+1}\|^{2}+\frac{c_{k-1}-c_{k}}{2}\|y_{k+1}\|^{2}.$ (59)
Proof  The optimality condition for $y_{k}$ in (43) implies that
$\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\gamma_{k}(y_{k+1}-y_{k}),y_{k}-y_{k+1}\rangle\leq 0,$ (60)
By (48) and (60), we have
$\langle\nabla_{y}f_{k-1}(x_{k},y_{k-1})-\nabla_{y}f_{k}(x_{k+1},y_{k})+$ $\gamma_{k}(y_{k+1}-y_{k})-\gamma_{k-1}(y_{k}-y_{k-1}),y_{k+1}-y_{k}\rangle\leq 0,$
and
$\gamma_{k}\|y_{k+1}-y_{k}\|^{2}-\gamma_{k-1}\langle y_{k+1}-y_{k},y_{k}-y_{k-1}\rangle$
$\leq$ $\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle,$ (61)
which can be rewritten as
$\left(\gamma_{k}-\frac{\gamma_{k-1}}{2}\right)\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}-\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$
$\leq$ $\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle$
$=$ $\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),v_{k+1}\rangle$
$+\langle\nabla_{y}f_{k}(x_{k-1},y_{k-1})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle.$ (62)
Using an argument similar to the proof of (50)-(53), by (C5) and the Cauchy-Schwarz inequality, we conclude from the above inequality that
$\left(\gamma_{k}-\frac{\gamma_{k-1}}{2}\right)\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}-\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$
$\leq$ $\;\frac{(l_{12}^{k})^{2}}{2a_{k}}\|x_{k+1}-x_{k}\|^{2}+\frac{a_{k}}{2}\|y_{k+1}-y_{k}\|^{2}-\frac{c_{k}-c_{k-1}}{2}(\|y_{k+1}\|^{2}-\|y_{k}\|^{2})$
$+\frac{c_{k}-c_{k-1}}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{1}{2\gamma_{k-1}}\|\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1})\|^{2}$
$+\frac{\gamma_{k-1}}{2}\|v_{k+1}\|-\frac{1}{{l}_{22}^{k-1}+2c_{k-1}}\|\nabla_{y}f_{k}(x_{k},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1})\|^{2}$
$-\frac{c_{k-1}{l}_{22}^{k-1}+c^{2}_{k-1}}{{l}_{22}^{k-1}+2c_{k-1}}\|y_{k}-y_{k-1}\|^{2},$ (63)
for any $a_{k}>0$. Obviously, we have $-\frac{c_{k-1}{l}^{k-1}_{22}+c^{2}_{k-1}}{2c_{k-1}+{l}^{k-1}_{22}}\leq-\frac{c_{k-1}{l}^{k-1}_{22}+c^{2}_{k-1}}{2({l}^{k-1}_{22}+c_{k-1})}=-\frac{c_{k-1}}{2}.$ Combining $\gamma_{k}\geq\frac{{l}^{k}_{22}+2c_{k}}{2}$ and rearranging the terms in (63), we obtain
$\frac{\gamma_{k}}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{c_{k}-c_{k-1}}{2}\|y_{k+1}\|^{2}$
$\quad\leq\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}+\left(\frac{\alpha_{k}}{2}+\frac{\gamma_{k-1}}{2}-\frac{\gamma_{k}}{2}\right)\|y_{k+1}-y_{k}\|^{2}+\frac{c_{k}-c_{k-1}}{2}\|y_{k}\|^{2}$
$\quad\quad+\frac{(l^{k}_{12})^{2}}{2a_{k}}\|x_{k+1}-x_{k}\|^{2}-\frac{c_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}.$ (64)
By multiplying $\frac{16\gamma_{k-1}}{c_{k-1}}$ on both sides of the above inequality, setting $a_{k}=\frac{c_{k}}{2}$, and using the definition of $S_{k+1}$, (58) and $\{\gamma_{k}\}$ is nodecreasing, we then obtain
$S_{k+1}-S_{k}$ $\leq 8\left({\gamma_{k}}\frac{c_{k+1}}{c_{k}}-{\gamma_{k-1}}\frac{c_{k}}{c_{k-1}}\right)\|y_{k+1}\|^{2}+\frac{16(l^{k}_{12})^{2}\gamma_{k-1}}{c_{k-1}^{2}}\|x_{k+1}-x_{k}\|^{2}$
$\quad+\frac{28\gamma_{k}}{5}\|y_{k+1}-y_{k}\|^{2}-{8}{\gamma_{k-1}}\|y_{k}-y_{k-1}\|^{2}.$ (65)
Combining (47) and (65), and using the definition of $F_{k+1}$, we complete the proof.
We are now ready to establish the iteration complexity for the PF-AGP-NC algorithm to achieve an $\varepsilon$-stationary point under the general nonconvex–concave setting.
Theorem 10
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 2. If $\beta_{k}=\frac{l_{12}^{k}}{20l^{k-1}_{22}}+\frac{1280(l_{12}^{k})^{2}\sqrt{k}}{19^{2}l_{12}^{k-1}}$, $\gamma_{k}=20l^{k}_{22}$ , $c_{k}=\frac{19l^{k}_{22}}{k^{1/4}}$, then for any given $\varepsilon>0$, $k\geq 2$, we have
$T(\varepsilon)\leq\max\left(\left(\frac{5120\left(L_{12}\right)^{2}d_{2}d_{3}}{19^{2}l_{22}^{1}\varepsilon^{2}}+1\right)^{2},\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}\right),$
where $d_{2}=F_{2}-\underline{F}+339L_{22}\hat{\mathcal{D}}_{y}^{2}$, $d_{3}=\max\{\bar{d}_{1},\frac{20\cdot 19^{2}L_{22}^{2}}{17(l_{12}^{1})^{2}}\}$, $\hat{\mathcal{D}}_{y}=\max\{\|y\|\,|y\in Y\}$ with $\underline{F}={\underline{f}}-639L_{22}\hat{\mathcal{D}}_{y}^{2}$, $\underline{f}=\min_{(x,y)\in X\times Y}f(x,y)$, $\bar{d}_{1}=32+\frac{1+1600L^{2}_{22}}{128({l}_{12}^{1})^{2}}$.
Proof  We can easily see that the relations in (58) are satisfied by the settings of $c_{k}$, $\gamma_{k}$ and $\beta_{k}$. Denote $\alpha_{k}=\frac{16(l_{12}^{k})^{2}\gamma_{k-1}}{c_{k-1}^{2}}$. Then it follows from the selection of $\beta_{k}$ and $\alpha_{k}$ that
$\frac{\beta_{k}}{2}-\frac{(l^{k}_{12})^{2}}{2\gamma_{k-1}}-\frac{16(l^{k}_{12})^{2}\gamma_{k-1}}{c_{k-1}^{2}}=\alpha_{k}.$ (66)
This observation, in view of Lemma 8, $\frac{c_{k}}{2}\leq\frac{19\gamma_{k}}{40k^{1/4}}\leq\frac{19\gamma_{k}}{40}$ implies that
$\begin{split}&{\alpha_{k}}\left\|x_{k+1}-x_{k}\right\|^{2}+\frac{17\gamma_{k}}{40}\left\|y_{k+1}-y_{k}\right\|^{2}\\ \leq&F_{k}-F_{k+1}+8\left({\gamma_{k}}\frac{c_{k+1}}{c_{k}}-{\gamma_{k-1}}\frac{c_{k}}{c_{k-1}}\right)\|y_{k+1}\|^{2}+\frac{c_{k-1}-c_{k}}{2}\|y_{k+1}\|^{2}.\end{split}$ (67)
We can easily check from the definition of $f_{k}(x,y)$ that
$\|\nabla G_{k}\|-\|\tilde{\nabla}G_{k}\|\leq c_{k}\|y_{k}\|.$ (68)
By replacing $f$ with $f_{k}$, similar to (32) and (33), we immediately obtain that
(~$\|(\tilde{\nabla}G_{k})_{x}\|=\beta_{k}\|x_{k+1}-x_{k}\|,$ (69)
and
(~$\|(\tilde{\nabla}G_{k})_{y}\|\leq\gamma_{k}\|y_{k+1}-y_{k}\|+l^{k}_{12}\|x_{k+1}-x_{k}\|.$ (70)
Combining (69) and (70), and using the Cauchy-Schwarz inequality, we have
~$\|\tilde{\nabla}G_{k}\|^{2}\leq(\beta_{k}^{2}+2(l^{k}_{12})^{2})\|x_{k+1}-x_{k}\|^{2}+2{{\gamma_{k}}^{2}}\|y_{k+1}-y_{k}\|^{2}.$ (71)
Since both $\alpha_{k}$ and $\beta_{k}$ are in the same order when $k$ becomes large enough, it then follows from the definition of $\bar{d}_{1}$ that $\forall k\geq 1$,
$\frac{(\beta_{k})^{2}+2({l^{k}_{12}})^{2}}{\alpha_{k}^{2}}$ $=\frac{\left(\frac{{l^{k}_{12}}}{\gamma_{k-1}}+\frac{64({l}_{12}^{k})^{2}\gamma_{k-1}}{c_{k-1}^{2}}\right)^{2}+2({l}^{k}_{12})^{2}}{\alpha_{k}^{2}}$
$\leq 32+\frac{1+1600L^{2}_{22}}{128({l}_{12}^{1})^{2}}=\bar{d}_{1}.$ (72)
Combining the previous two inequalities in (71) and (72), we obtain
$\|\nabla{\tilde{G}_{k}}\|^{2}\leq{\bar{d}_{1}(\alpha_{k})^{2}}\|x_{k+1}-x_{k}\|^{2}+2\gamma_{k}^{2}\|y_{k+1}-y_{k}\|^{2}.$ (73)
Denote $d_{k}^{(2)}=\frac{1}{\max\{\bar{d}_{1}\alpha_{k},{\frac{17}{80}\gamma_{k}}\}}$, $\tilde{c}_{k}=\frac{19}{20k^{1/4}}$, by multiplying $d^{(2)}_{k}$ on the both sides of (73), and using (67), we have
$d_{k}^{(2)}\|\nabla\tilde{G}_{k}\|^{2}$
$\leq$ $F_{k}-F_{k+1}+8\left({\gamma_{k}}\frac{c_{k+1}}{c_{k}}-{\gamma_{k-1}}\frac{c_{k}}{c_{k-1}}\right)\|y_{k+1}\|^{2}+\frac{c_{k-1}-c_{k}}{2}\|y_{k+1}\|^{2}$
$\leq$ $F_{k}-F_{k+1}+8\left({\gamma_{k}}\frac{c_{k+1}}{c_{k}}-{\gamma_{k-1}}\frac{c_{k}}{c_{k-1}}\right)\|y_{k+1}\|^{2}+40L_{22}\frac{\tilde{c}_{k-1}-\tilde{c}_{k}}{2}\|y_{k+1}\|^{2}.$ (74)
Denoting
$\tilde{T}(\varepsilon)=\min\left\{k\middle|\left\|\nabla\tilde{G}(x_{k},y_{k})\right\|\leq\frac{\varepsilon}{2},k\geq 2\right\},$ (75)
summing both sides of (74) from $k=2$ to $k=\tilde{T}(\varepsilon)$, we then obtain
$\sum_{k=2}^{\tilde{T}{(\varepsilon)}}d^{(2)}_{k}\left\|\nabla\tilde{G}_{k}\right\|^{2}$ $\leq F_{2}-F_{\tilde{T}(\varepsilon)}+8\left(\frac{\gamma_{\tilde{T}(\varepsilon)-1}c_{\tilde{T}(\varepsilon)}}{c_{\tilde{T}(\varepsilon)-1}}-\frac{\gamma_{1}c_{2}}{c_{1}}\right)\hat{\mathcal{D}}_{y}^{2}+40L_{22}\left(\frac{\tilde{c}_{1}}{2}-\frac{\tilde{c}_{\tilde{T}(\varepsilon)}}{2}\right)\hat{\mathcal{D}}^{2}_{y}$
$\leq F_{2}-F_{\tilde{T}(\varepsilon)}+339L_{22}\hat{\mathcal{D}}_{y}^{2}.$ (76)
Note that by the definition of $F_{k+1}$ in Lemma 9, we have
$\begin{split}F_{\tilde{T}(\varepsilon)}&\geq{\underline{f}}-639L_{22}\hat{\mathcal{D}}_{y}^{2}=\underline{F},\end{split}$ (77)
where $\underline{f}=\min_{(x,y)\in X\times Y}f(x,y)$. We then conclude from (76) that
$\sum_{k=2}^{\infty}d^{(2)}_{k}\left\|\nabla\tilde{G}_{k}\right\|^{2}\leq F_{2}-\underline{F}+339L_{22}\hat{\mathcal{D}}_{y}^{2}=d_{2}.$ (78)
We can see from the selection of $d_{3}$ that $d_{3}\geq\max\{\bar{d}_{1},\frac{80\gamma_{k}}{17\alpha_{k}}\}$, which implies $d^{(2)}_{k}\geq\frac{1}{d_{3}\alpha_{k}}$, by multiplying $d_{3}$ on the both sides of (78), and combining the definition of $d_{2}$, we have $\sum_{k=2}^{\infty}\frac{1}{\alpha_{k}}\left\|\nabla\tilde{G}_{k}\right\|^{2}\leq d_{2}d_{3}$, which, by the definition of $\tilde{T}(\varepsilon)$, implies that
$\frac{\varepsilon^{2}}{4}\leq\frac{d_{2}d_{3}}{\sum_{k=2}^{\tilde{T}{(\varepsilon)}}\frac{1}{\alpha_{k}}}\leq\frac{d_{2}d_{3}}{\sum_{k=2}^{\tilde{T}{(\varepsilon)}}\frac{1}{\bar{\alpha}}}.$ (79)
By using the fact $\sum_{k=2}^{\tilde{T}{(\varepsilon)}}\frac{1}{\sqrt{k}}\geq\sqrt{\tilde{T}{(\varepsilon)}}-1$ and (79), we conclude $\frac{\varepsilon^{2}}{4}\leq\frac{1280\cdot\left(L_{12}\right)^{2}d_{2}d_{3}}{19^{2}l_{22}^{1}\left(\sqrt{\tilde{T}(\varepsilon)}-1\right)}$ or equivalently,
$\tilde{T}(\varepsilon)\leq\left(\frac{5120\left(L_{12}\right)^{2}d_{2}d_{3}}{19^{2}l_{22}^{1}\varepsilon^{2}}+1\right)^{2}.$ (80)
On the other hand, if $k\geq\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}$, then $c_{k}=\frac{19\gamma_{k}}{20k^{1/4}}\leq\frac{\varepsilon}{2\hat{{\mathcal{D}_{y}}}}$. This inequality together with the definition of $\hat{\mathcal{D}}_{y}$, then imply that $c_{k}\left\|y_{k}\right\|\leq\frac{\varepsilon}{2}$. Therefore, there exists a
$T(\varepsilon)$ $\leq\max\left(\tilde{T}{(\varepsilon)},\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}\right)$
$\leq\max\left(\left(\frac{5120\left(L_{12}\right)^{2}d_{2}d_{3}}{19^{2}l_{22}^{1}\varepsilon^{2}}+1\right)^{2},\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}\right),$ (81)
such that $\left\|\nabla G_{k}\right\|\leq\left\|\nabla\tilde{G}_{k}\right\|+c_{k}\left\|y_{k}\right\|\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon$.
Remark 11
At the $k$-th iteration of Algorithm 2, the number of backtracking step is upper bounded by $\log_{2}\left(\frac{2L_{11}}{l_{11}^{1}}\right)+\log_{2}\left(\frac{2L_{12}}{l_{12}^{1}}\right)+\log_{2}\left(\frac{2L_{22}}{l_{22}^{1}}\right)=\mathcal{O}(1)$. According to Theorem 10, the total number of gradient calls of Algorithm 2 to obtain an $\varepsilon$-stationary point that satisfies $\|\nabla G(x_{k},y_{k})\|\leq\varepsilon$ is $\mathcal{O}(L^{4}\varepsilon^{-4})$ under the nonconvex-concave setting.
3 Numerical results
In this section, we report some numerical results to show the efficiency of the proposed PF-AGP-NSC and PF-AGP-NC algorithms. All the numerical tests are implemented in Python 3.12 and run with an NVIDIA RTX 4090D GPU (24GB).
3.1 Synthetic Minimax Problem
Synthetic minimax problem can be formulated as the following nonconvex-strongly-concave minimax problem (Luo et al., 2022):
$\min_{\mathbf{x}\in\mathbb{R}^{3}}\max_{\mathbf{y}\in\mathbb{R}^{2}}f(\mathbf{x},\mathbf{y})=w(x_{3})-\frac{y_{1}^{2}}{40}+x_{1}y_{1}-\frac{5y_{2}^{2}}{2}+x_{2}y_{2},$ (82)
where $\mathbf{x}=[x_{1},x_{2},x_{3}]^{\top}$, $\mathbf{y}=[y_{1},y_{2}]^{\top}$ and
$w(x)=\begin{cases}\sqrt{\epsilon}\left(x+(\lambda+1)\sqrt{\epsilon}\right)^{2}-\frac{1}{3}\left(x+(\lambda+1)\sqrt{\epsilon}\right)^{3}-\frac{1}{3}(3\lambda+1)\epsilon^{3/2},&x\leq-\lambda\sqrt{\epsilon};\\ \epsilon x+\frac{\epsilon^{3/2}}{3},&-\lambda\sqrt{\epsilon}<x\leq-\sqrt{\epsilon};\\ -\sqrt{\epsilon}x^{2}-\frac{x^{3}}{3},&-\sqrt{\epsilon}<x\leq 0;\\ -\sqrt{\epsilon}x^{2}+\frac{x^{3}}{3},&0<x\leq\sqrt{\epsilon};\\ -\epsilon x+\frac{\epsilon^{3/2}}{3},&\sqrt{\epsilon}<x\leq\lambda\sqrt{\epsilon};\\ \sqrt{\epsilon}\left(x-(\lambda+1)\sqrt{\epsilon}\right)^{2}+\frac{1}{3}\left(x-(\lambda+1)\sqrt{\epsilon}\right)^{3}-\frac{1}{3}(3\lambda+1)\epsilon^{3/2},&\lambda\sqrt{\epsilon}\leq x.\end{cases}$
It is easy to verify that this problem has a strict saddle point at $(\mathbf{x}^{*},\mathbf{y}^{*})=([0,0,0]^{\top},[0,0]^{\top})$.
Experimental setup. We compare the proposed PF-AGP-NSC algorithm with the SGDA-B algorithm, the TiAda algorithm and the AGP algorithm for solving (82) with $\epsilon=0.01$, $\lambda=5$. The initial point is chosen as $(x_{0},y_{0})=([0,0,2]^{\top},[0,0]^{\top})$ for all four tested algorithms. For the PF-AGP-NSC algorithm, we set $l_{11}^{1}=0.01$ , $l_{12}^{1}=0.01$, $l_{22}^{1}=0.01$ and $\mu_{1}=1$. For the SGDA-B algorithm, we set $\mu=1$, $\gamma=0.9$, $N=1$, $p=0.5$ and $\tilde{L}=\frac{10}{9}$. For the TiAda algorithm, we set $\alpha=0.6$, $\beta=0.4$, $\eta_{x}=0.8$, $\eta_{y}=0.4$, $v^{x}_{0}=v^{y}_{0}=1$. For the AGP algorithm, we set $\alpha_{x}=0.14$, $\beta_{y}=1.1$. The algorithms terminate when $\|\nabla f(x_{k},y_{k})\|\leq 10^{-5}$ satisfied.
Figure 1: Numerical Performance of four tested algorithms for solving the Synthetic Minimax problem
Results. Figure 1 shows the stationarity gap of all four test algorithms, which demonstrates that the proposed PF-AGP-NSC algorithm exhibits significantly better convergence performance than that of the SGDA-B algorithm, TiAda algorithm and AGP algorithm.
3.2 Dirac-GAN problem
The Dirac-GAN problem can be formulated as the following nonconvex-concave minimax problem (Mescheder et al., 2018):
$\min_{x}\max_{y}L(x,y)=-\log(1+\exp(-xy))+\log 2,$ (83)
where $(0,0)$ is the unique stationary point.
Experimental setup. We compare the PF-AGP-NC algorithm with the SGDA-B algorithm and the AGP algorithm. The initial point of all algorithms is chosen as $(1,1)$. For the PF-AGP-NC algorithm, we set $l_{11}^{1}=0.01$ , $l_{12}^{1}=1$ and $l_{22}^{1}=0.01$. For the SGDA-B algorithm, we set $\mu=1$, $\gamma=0.5$, $N=1$, $p=0.5$ and $\tilde{L}$ = 1. For the AGP algorithm, we set $\alpha_{x}=\frac{0.8}{\sqrt{k}}$, $\beta_{y}=0.3$ and $c_{k}=\frac{0.5}{k^{1/4}}$. The algorithms terminate when $\|\nabla f(x_{k},y_{k})\|\leq 10^{-5}$ satisfied.
Results. Figure 2 shows the stationarity gap and the sequences generated by all three test algorithms. We find that the SGDA-B algorithm fails to converge, both the AGP algorithm and the PF-AGP-NC algorithms successfully approach the unique stationary point. Moreover, the PF-AGP-NC algorithm exhibits significantly better convergence performance than that of the AGP algorithm.
Figure 2: Numerical performance of three tested algorithms for solving the Dirac-GAN problem
3.3 Robust learning over multiple domains
In this subsection, we consider the following robust learning problem over multiple domains (Qian et al., 2019), formulated as a nonconvex-linear minimax problem:
$\min_{x}\max_{y\in\Delta}y^{T}F(x),$ (84)
where $F(x):=\left[f_{1}(x);\ldots;f_{M}(x)\right]\in\mathbb{R}^{M\times 1}$ with $f_{m}(x)=\frac{1}{\left|\mathcal{S}_{m}\right|}\sum_{i=1}^{\left|\mathcal{S}_{m}\right|}\ell\left(d_{i}^{m},l_{i}^{m},x\right)$ and $\ell$ can be any non-negative loss function, $M$ is the number of tasks and $x$ denotes the network parameters, $y\in\Delta$ describes the weights over different tasks and $\Delta$ is the simplex, i.e., $\Delta=\left\{(y_{1},\cdots,y_{M})^{T}\mid\sum_{m=1}^{M}y_{m}=1,y_{m}\geq 0\right\}$. We consider two image classification problems with MNIST (LeCun et al., 1998) and CIFAR10 datasets (Krizhevsky et al., 2009). Our goal is to train a neural network that works on these two completely unrelated problems simultaneously. Since cross-entropy loss is popular in multi-class classification problem, we use it as our loss function. We should point out that the quality of the algorithm for solving a robust learning problem over multiple domains is measured by the worst case accuracy over all domains (Qian et al., 2019).
Experiment setup. We compare the proposed PF-AGP-NC algorithm with the SGDA-B algorithm and the AGP algorithm. For the PF-AGP-NC algorithm, we set $l_{11}^{1}=0.01$, $l_{12}^{1}=0.01$ and $l_{22}^{1}=0.01$. For the SGDA-B algorithm, we set $\mu=20$, $\gamma=0.9$, $N=1$, $p=0.5$ and $\tilde{L}$ = 20. For the AGP algorithm, we set $\frac{1}{\beta_{k}}=\frac{2}{2+\sqrt{k}}$, $\gamma_{k}=100$ and $c_{k}=\frac{1}{10+k^{\frac{1}{4}}}$. We set the batch size in all tasks to be $128$ and run $50$ epochs for all algorithms. Moreover, we sample our results every epoch and calculate the average accuracy on these two testing sets to evaluate the performance of different algorithms. Due to hardware (especially memory) limitations, we use a mini-batch randomized gradient instead of the exact gradient at each iteration in our numerical experiment, which is the same as in (Xu et al., 2023).
Figure 3: Performance of two algorithms for solving robust multi-task learning problem.
Results. Figure 3 and Figure 3 show the testing accuracy on the MNIST dataset and the CIFAR10 dataset respectively. PF-AGP-NC algorithm outperforms the SGDA-B algorithm and the AGP algorithm for both datasets.
4 Conclusions
In this paper, based on the framework of backtracking, we propose two completely parameter-free alternating gradient projection algorithms, i.e., the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm, for solving nonconvex-(strongly) concave minimax problems respectively, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concave constant $\mu$. Moreover, we show that the total number of gradient calls of the PF-AGP-NSC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-strongly concave minimax problems is upper bounded by $\mathcal{O}\left(L\kappa^{3}\varepsilon^{-2}\right)$ where $\kappa$ is the condition number, while the total number of gradient calls of the PF-AGP-NC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-concave minimax problems is upper bounded by $\mathcal{O}\left(L^{4}\varepsilon^{-4}\right)$. As far as we know, the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm are the first completely parameter-free algorithms for solving nonconvex-strongly concave minimax problems and nonconvex-concave minimax problems respectively. Numerical results demonstrate the efficiency of the two proposed algorithms.